OpenStudy (anonymous):

Write all of the following statement(s) that apply to the following polynomial: -5x^3 + 2x^2 – 7 It is a binomial. It cannot be evaluated at x = -1/3 It has a degree of 3. It is cubic.

OpenStudy (anonymous):

It is a binomial I guess.

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

Yes that's it, anything else?

OpenStudy (anonymous):

so just say its that

OpenStudy (anonymous):

Lolz wow

OpenStudy (istim):

I think it's everything but I haven't checked "It cannot be evaluated at x = -1/3". IF you wanna check ,sub that in the equation, and see if the result is good.

OpenStudy (anonymous):

I am lost

OpenStudy (istim):

I don't think its a binomial, but what do I know.

OpenStudy (istim):

"It is a binomial." Forgot what binomial is... "It cannot be evaluated at x = -1/3" Sub x=-1/3 into -5x^3 + 2x^2 – 7. It has a degree of 3.- I think of -5x^3. "It is cubic." I think of "-5x^3 ".

OpenStudy (anonymous):

I been told you it's a binomial. Wow.

OpenStudy (istim):

Looks to me as a trinomial, or greater.

OpenStudy (anonymous):

It's a binomial dude.

OpenStudy (istim):

What's its binomial form? (x-a)(x-b) or whatever? It looks to be that right?

OpenStudy (istim):

Sorry, getting off what the questioner wants, but I really should learn this.

OpenStudy (anonymous):

Than Tim don't confuse her man. It's a binomial for the 100 th time I said it. lol

OpenStudy (anonymous):

u already confused me

OpenStudy (bahrom7893):

ITS NOT BINOMIAL!

OpenStudy (bahrom7893):

HELLOOO BI means two! It's cubic and has a degree of 3, it's defined everywhere, so undefined at 1/3 is false too

OpenStudy (bahrom7893):

C and D are the options

OpenStudy (anonymous):

so the third and fourth one are correct

OpenStudy (istim):

You gotta sub in the value to check the second one. It's very possible for it to work.

OpenStudy (istim):

I'm not doing it myself cause I'm working on something else.

OpenStudy (bahrom7893):

Listen the only way it wouldnt work if u had a denominator with x + 1/3 cuz u'd get a 0 on the denominator.. or something like a negative under sqrt. YES THE ANSWER IS 3 AND 4.

OpenStudy (istim):

Ok. Makes sense.

OpenStudy (istim):

Didn't check it out personally (mathematically).